OpenStudy (anonymous):

1500=200e^(25r) solve for r

OpenStudy (unklerhaukus):

divide by 200, then take the natural log of both sides

OpenStudy (unklerhaukus):

then divide by 25

OpenStudy (anonymous):

so log(1500/200)=25r log(e)?

OpenStudy (anonymous):

is that right?

OpenStudy (unklerhaukus):

natural log of x is \[\log_e(x)=\ln(x)\]

OpenStudy (unklerhaukus):

so \[\log_e(e)=\ln (e)=1\]

OpenStudy (anonymous):

so log (e) =1?

OpenStudy (anonymous):

i dint get it

OpenStudy (anonymous):

mind you e is the mathematical constant. i.e. e=2.718....

OpenStudy (unklerhaukus):

when you write log it means the log to base 10\[\log (x)=\log_{10}(x)\] so \[\log (10)=\log_{10}(10)=1\] ___ but in theis question the base is e, so take the natural log instead of the log to base ten

OpenStudy (anonymous):

so is hould ln(e)

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

so \[\ln(1500/200)=25r \ln(e)\]

OpenStudy (unklerhaukus):

now the ln (e) but simplifies to one

OpenStudy (anonymous):

ln(1500/200)= 2.105 ln(1500/200)/ln(e)=2.105?

OpenStudy (unklerhaukus):

/?

OpenStudy (anonymous):

i put ln(1500/200) into the calculator. it gave me 2.105

OpenStudy (anonymous):

then i divided it by ln(e)

OpenStudy (unklerhaukus):

oh, i wouldn't make an approximation \[\ln(1500/200)=\ln(15/2)=\ln(7.5)\]or\[\ln(1500/200)=\ln(15/2)=\ln(15)-\ln(2)\]

OpenStudy (anonymous):

yes. and ln(7.5) = 2.104903021

OpenStudy (unklerhaukus):

yeah, if you were going to approximate it

OpenStudy (anonymous):

ok so ln(7.5)/ln(e)=ln(7.5)

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

so ln(7.5)/25 is the answer?

OpenStudy (anonymous):

= .0805961208

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

yeah that is a rough value for r ,

OpenStudy (unklerhaukus):

if i was you i would leave my answer as r=ln(7.5)/25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see the real question was The count in a bateria culture was 200 after 15 minutes and 1500 after 40 minutes. What was the initial size of the culture? Find the doubling period. Find the population after 120 minutes. When will the population reach 11000.

OpenStudy (anonymous):

i got everything but population after 120 mins.

Parth (parthkohli):

UnkleRhaukus is right. Also,\[\rm \large \ln|e^{blah}|= blah\]

OpenStudy (unklerhaukus):

a general equation for growth is\[A(t)=A_0 e^{t/\tau}\] where A(t) is the amount at time t A_0 is the initial amount tau is the time for the amount to increase by a factor of e \[1500(r)=200e^{25r} \]

OpenStudy (anonymous):

ok... so my formula was wrong

OpenStudy (unklerhaukus):

your formula was fine

OpenStudy (unklerhaukus):

i think i made a typo though

OpenStudy (unklerhaukus):

\[A(r)=200e^{25r}\] to find the amount at r=120 plug in r=120into the RHS

OpenStudy (unklerhaukus):

i have to go now ,

OpenStudy (anonymous):

so a(r) = 200e^(25*120)>

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

yeah thats it , just simplify (using a calculator), and round down to the nearest whole number to find A at r=120

OpenStudy (anonymous):

calculator cannot calculate bcoz of overflow